The catalysts decrease the Ea from 100 K J mol^-1 to 80 K J mol^-1 . At what temperature the rate of reaction in the absence of catalyst at 500 K will be equal to rate of reaction in presence of catalyst .

Grade:11

2 Answers

Mujahid Ahmed
11 years ago

Apply Arhenius equation.

K = Ae^-Ea/RT

 

If the two Rates are equal. 

e^-Ea1/RT1 = e^-Ea2/RT2

Ea1/T1 = Ea2/T2

 

100/T = 80/500

T = 625 K

 

Approve if i helped.

ashish chauhan
6 years ago
the catalyst decreses the Ea from 100 kj mol to 80 kj mol .at what temperature the rate of reaction in absence of catalyst at 500 k will be equal to rate of reaction in presence of catalyst